Advertisements
Advertisements
प्रश्न
The time to failure in thousands of hours of an electronic equipment used in a manufactured computer has the density function
`f(x) = {{:(3"e"^(-3x), x > 0),(0, "eleswhere"):}`
Find the expected life of this electronic equipment
Advertisements
उत्तर
`mu = "E"("X")`
= `int_0^oo x f(x) "d"x`
= `3 int_0^oo x"e"^(-3x) "d"x`
u = x, `int"dv" = int"e"^(-3x)"d"x`
du = dx, v = `"e"^(-3x)/(-3)`
`int"udv" = "uv" - int"vdu"`
= `3[(x"e"^(-3x))/(-3) + int"e"^(-3x)/3 "d"x]_0^oo`
= `3[- (x"e"^(-3x))/3 - "e"^(-3x)/9]_0^oo`
= `3[0 - (0 - 1/9)]`
= `1/3`
Expected life of the electronic equipment = `1/3`
APPEARS IN
संबंधित प्रश्न
For the random variable X with the given probability mass function as below, find the mean and variance.
`f(x) = {{:(1/10, x = 2"," 5),(1/5, x = 0"," 1"," 3"," 4):}`
For the random variable X with the given probability mass function as below, find the mean and variance.
`f(x) = {{:(1/2 "e"^(x/2), "for" x > 0),(0, "otherwise"):}`
Two balls are drawn in succession without replacement from an urn containing four red balls and three black balls. Let X be the possible outcomes drawing red balls. Find the probability mass function and mean for X
A lottery with 600 tickets gives one prize of ₹ 200, four prizes of ₹ 100, and six prizes of ₹ 50. If the ticket costs is ₹ 2, find the expected winning amount of a ticket
Choose the correct alternative:
On a multiple-choice exam with 3 possible destructive for each of the 5 questions, the probability that a student will get 4 or more correct answers just by guessing is
Choose the correct alternative:
If P(X = 0) = 1 – P(X = 1). If E[X] = 3 Var(X), then P(X = 0) is
How do you defi ne variance in terms of Mathematical expectation?
In a business venture a man can make a profit of ₹ 2,000 with a probability of 0.4 or have a loss of ₹ 1,000 with a probability of 0.6. What is his expected, variance and standard deviation of profit?
Choose the correct alternative:
Value which is obtained by multiplying possible values of a random variable with a probability of occurrence and is equal to the weighted average is called
Choose the correct alternative:
Given E(X) = 5 and E(Y) = – 2, then E(X – Y) is
Choose the correct alternative:
A discrete probability distribution may be represented by
Choose the correct alternative:
A probability density function may be represented by
Choose the correct alternative:
E[X – E(X)] is equal to
Choose the correct alternative:
E[X – E(X)]2 is
Choose the correct alternative:
A listing of all the outcomes of an experiment and the probability associated with each outcome is called
Choose the correct alternative:
The distribution function F(x) is equal to
What is the expected value of a game that works as follows: I flip a coin and if tails pay you ₹ 2; if heads pay you ₹ 1. In either case, I also pay you ₹ 0.50
Prove that V(aX) = a2V(X)
Prove that V(X + b) = V(X)
